OpenStudy (anonymous):

can anyone help me on what are the rules in factoring both perfect square??thanks

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

so if you see \[x^2+10x+25\] you see at a = x, b = 5 and you write \[(x+5)^2\]

OpenStudy (anonymous):

\[x^2+7x+\frac{49}{4}\] you thing a = x, \[b=\frac{7}{2}\] and write \[(x+\frac{7}{2})^2\]

OpenStudy (anonymous):

unless you meant \[a^2-b^2=(a+b)(a-b)\] which is called "the DIFFERENCE of two squares"

OpenStudy (anonymous):

so for example \[x^2-100\] you can rewrite as \[(x+10)(x-10)\]